Find the radiation resistance of an antenna of length λ/10 meter?


1
Expert's answer
2021-01-06T13:53:06-0500
"R=80\pi^2\left(\frac{dl}{\lambda}\right)^2\\
R=80\pi^2\left(\frac{\lambda}{10\lambda}\right)^2=7.9\Omega"